### v1

@@ Abstract
We present ActionDiffusion – a novel diffusion model for procedure planning in instructional videos that is the first to take temporal inter-dependencies between actions into account in a diffusion model for procedure planning. This approach is in stark contrast to existing methods that fail to exploit the rich information content available in the particular order in which actions are performed. Our method unifies the learning of temporal dependencies between actions and denoising of the action plan in the diffusion process by projecting the action information into the noise space. This is achieved 1) by adding action embeddings in the noise masks in the noise-adding phase and 2) by introducing an attention mechanism in the noise prediction network to learn the correlations between different action steps. We report extensive experiments on three instructional video benchmark datasets (CrossTask, Coin, and NIV) and show that our method outperforms previous state-of-the-art methods on all metrics on CrossTask and NIV and all metrics except accuracy on Coin dataset. We show that by adding action embeddings into the noise mask the diffusion model can better learn action temporal dependencies and increase the performances on procedure planning.
Index Terms:
Procedure planning, diffusion models, temporal dependency
I Introduction
@@
To support humans in everyday procedural tasks, such as cooking or cleaning, future autonomous robots need to be able to plan actions from visual observations of human actions and their environment – so-called procedure planning [1, 2]. Procedure planning is commonly defined as the task of predicting an action plan, i.e., a sequence of individual actions, from only a start and final observation of the overall procedure. Several previous works in robotics have investigated procedure planning from visual observations [3, 4, 5]. But these works have learnt visual representations from artificial and rather simple images, such as a simulated cart pole. Other works have used real-world images to learn action plans [6, 7] but the environment was simplified and constrained by pre-defined object-centred representations [8], for example, coloured cubes as objects on a table.
Latest works in robotics still struggle with learning plans from the natural demonstration of tasks by humans [9, 10, 11] since the human demonstrations are performed in complex, unstructured natural environments. Leveraging more advanced deep learning methods for planning procedures from instructional videos has the potential to address this limitation. In [2], the authors have used convolutional neural networks (CNNs) to extract visual features and modelled the dynamics between observations and actions in the feature space using multi-layer perceptrons (MLPs) and recurrent neural networks (RNNs). In [7], the authors followed the same paradigm of modelling the dynamics between actions and observations but used transformers instead. RL was also used in [12] to learn the action policy. Common to all of these works is their use of a separate planning algorithm (beam search or walk-through) during planning. Later works used generative models to sample action plans during the inference stage [12, 13]. Although different approaches have been developed to tackle the procedure planning task, the challenge remains open due to the complex and unstructured video observations.
@@
Figure 1: Procedure planning in instructional videos using diffusion models. Previous work (Projected Diffusion) [14] does not take the temporal dependencies between actions into account. Our method (ActionDiffusion) incorporates these dependencies directly into the diffusion model.
@@
Diffusion models [15] have achieved outstanding results in many research fields such as in image generation [15, 16], text-to-image generation [17, 18, 19], trajectory planning [20, 21], video generation [22, 23], human motion prediction [24], of time series imputation [25]. The current state-of-the-art method for procedure planning in videos [14] is also based on a diffusion model. Unlike diffusion models for images, the input for diffusion-based procedure planning is a multi-dimensional matrix that consists of the visual observations of the start and goal, the sequence of actions, and task classes. At inference time, the input matrix is reconstructed and the dimension containing the action sequence is the prediction of the action plan. A key limitation of this method is that the influence of the temporal dependencies between actions, i.e. that actions are more likely to cause particular follow-up actions, has not been considered. Although the input matrix in [14] contains the action labels during training, the diffusion model still treats the input matrix as a static “image”, it does not learn its inherent temporal dependencies.
@@
To overcome this limitation we propose Action-aware Noise Mask Diffusion (ActionDiffusion) – the first method that leverages the temporal dependencies between actions into the diffusion process (see Figure 1). Our method learns the temporal dependencies of actions in the noise space of the diffusion instead of the feature space, thereby unifying the tasks of learning temporal dependencies and generating action plans in the diffusion steps. More specifically, we first propose an action-aware noise mask for the noise-adding stage of the diffusion model. We add the action embeddings in addition to the Gaussian noise in the noise masks so that the model input is transformed into Gaussian noises by iteratively adding the action-aware noise. Second, we introduce an attention mechanism in the denoising neural network (U-Net) to learn the correlations between actions. During the inference, it then can predict action-aware noises to generate action plans. In line with previous work [13, 14], we evaluate the performance of ActionDiffusion through experiments on three popular instructional video benchmarks: CrossTask [9], Coin [10], and NIV [11] with various time horizons. Our results show that ActionDiffusion achieves State-Of-The-Art (SOTA) performances across different metrics for all time horizons on all three datasets. We demonstrate that by adding action embeddings into the noise mask, the diffusion model can effectively learn the temporal dependencies between actions and increase the performance on procedure planning in instructional videos.
@@
II Method
@@
II-A Problem Formulation
We adopt the problem formulation of procedure planning used in previous work [2, 14]: Given the visual observation of the start state $o_{s}$ and the goal state $o_{g}$, the task is to predict the intermediate steps, i.e., the action plan $\pi=a_{1:T}$ for a chosen time horizon $T$. The action plan $\pi$ will transform $o_{s}$ to $o_{g}$. More formally, this task can be written as,
@@ II-A Problem Formulation
where $\hat{c}$ is the predicted task class of the video (e.g., cooking a specific meal). The planning is decomposed into two steps [14]: 1) predicting the task class $\hat{c}$ given the start state $o_{s}$ and the goal state $o_{g}$, 2) inferring the action plan $\pi$ given $o_{s}$, $o_{g}$, and $\hat{c}$ by sampling from the diffusion model.
II-B Action-aware Noise Mask Diffusion
Figure 2 provides an overview of our proposed ActionDiffusion method. Taking an instructional video as input, the method first extracts the visual features of the start state $o_{s}$, the goal state $o_{g}$, and of action embeddings $a_{e_{1:T}}$. It then uses the task class $c$, one-hot action class $a_{1:T}$, $o_{s}$, and $o_{g}$ to form the input for the diffusion model. Note that during training, we use the ground truth task class $c$ while the predicted task class $\hat{c}$ is used during inference. To obtain $\hat{c}$ we train a separate task classifier that models $p(\hat{c}\mid o_{s},o_{g})$ in Equation (1). In the noise-adding phase, during training, the noise is added on $a_{1:T}$ in the model input $x_{0}$. For each action, we add all previous as well as the current action embedding in addition to the Gaussian noise. In the denoising phase, during inference, we use a U-Net with attention to predict the action-aware noise to denoise the noisy input $x_{n}$ at step $n$. The action sequence $\hat{a}_{1:T}$ from the reconstructed input $x_{0}$ is the predicted action plan (see below for details).
II-C Diffusion Model
@@
II-D Action-aware Noise Mask
@@
at the noise-adding step, the noise is only added to the action dimension, i.e. $a_{1:T}$ and $a_{i}$ is a one-hot vector of a number of $A$ action classes, thus $\mathbf{I}\in\mathbb{R}^{T\times A}$ in Equation (2). We design an action-aware noise mask $M_{a}$ with multiple previous actions accumulated (multi-add) as
@@
where $a_{e_{1:T}}$ are the embeddings of $a_{1:T}$, and $g(\cdot)$ normalises the embeddings to the range of [-1, 1]. Essentially, in addition to the Gaussian noise applied to $x_{0}$, we add the normalised action embeddings to the noise mask. In the temporal direction, we accumulate all previous action embeddings. Our intuition is that at each $i\in[2:T]$ the noise mask knows what the previous actions are and $\epsilon_{\theta}$ can learn the temporal dependencies of actions in the denoising stage. With the action-aware noise mask, Equation (2) then becomes,
@@
we again apply re-parameterisation and update Equation (3),
$x_{n}=\sqrt{\bar{\alpha}_{n}}x_{0}+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{a}.$ (9)
II-E Denoising Neural Network
@@
Following [14], we use a U-Net as the noise prediction neural network $\epsilon_{\theta}$. To better learn the temporal dependencies between actions we further propose to incorporate an attention mechanism [26, 17]. Figure 3 shows the architecture of the U-Net with attention. As indicated in Equation (5), the U-Net takes $x_{n}$ and noise schedule time step $n$ as input. $n$ is first passed to a time step block and then is fed to all residual blocks. The attention block enables self-attention [27].
II-F Training
@@
$\mathcal{L}=\sum_{n=1}^{N}(\epsilon_{\theta}(x_{n},n),x_{0})^{2}.$ (10)
II-G Inference
During inference, only $o_{s}$ and $o_{g}$ are given and we need to sample the action plan $\pi$. We start with constructing $x_{n}$. Then, the noise prediction neural network $\epsilon_{\theta}$ predicts the noise iteratively to denoise $x_{n}$ into $x_{0}$. Here, $x_{n}$ is
$x_{n}=\begin{bmatrix}\hat{c}&\hat{c}&...&\hat{c}&\hat{c}\\ \epsilon&\epsilon&...&\epsilon&\epsilon\\ o_{s}&0&...&0&o_{g}\end{bmatrix},$ (11)
only the action plan is initialised with Gaussian noise and $\epsilon\in\mathcal{N}(0,\mathbf{I})$ since the noise is only added to the actions during the noise-adding stage. Although we use action-aware noise masks in the noise-adding stage, we do not need to encode action information into the Gaussian noise. The reason is that, after adding noise to $x_{0}$ with $N$ steps, $x_{N}$ already approximates Gaussian noise by construction. After $N$ steps of denoising, we take the action sequence $\hat{a}_{1:T}$ as the action plan $\pi$.
III Experiments
III-A Datasets
@@
III-B Metrics

### v2

@@ Abstract
We present ActionDiffusion – a novel diffusion model for procedure planning in instructional videos that is the first to take temporal inter-dependencies between actions into account. Our approach is in stark contrast to existing methods that fail to exploit the rich information content available in the particular order in which actions are performed. Our method unifies the learning of temporal dependencies between actions and denoising of the action plan in the diffusion process by projecting the action information into the noise space. This is achieved 1) by adding action embeddings in the noise masks in the noise-adding phase and 2) by introducing an attention mechanism in the noise prediction network to learn the correlations between different action steps. We report extensive experiments on three instructional video benchmark datasets (CrossTask, Coin, and NIV) and show that our method outperforms previous state-of-the-art methods on all metrics on CrossTask and NIV and all metrics except accuracy on Coin dataset. We show that by adding action embeddings into the noise mask the diffusion model can better learn action temporal dependencies and increase the performances on procedure planning.
1 Introduction
@@
To support humans in everyday procedural tasks, such as cooking or cleaning, future autonomous AI agents need to be able to plan actions from visual observations of human actions and their environment – so-called procedure planning [16, 3]. Procedure planning is commonly defined as the task of predicting an action plan, i.e., a sequence of individual actions, from only a start and final observation of the overall procedure. Several previous works have investigated procedure planning from visual observations [9, 27, 8]. But these works have learnt visual representations from artificial and rather simple images, such as a simulated cart pole. Other works have used real-world images to learn action plans [5, 24] but the environment was simplified and constrained by pre-defined object-centred representations [15], for example, coloured cubes as objects on a table. Leveraging more advanced deep learning methods for planning procedures from instructional videos has the potential to address this limitation [3, 2, 34]. Although different approaches have been developed to tackle the procedure planning task, the challenge remains open due to the complex and unstructured video observations.
@@
Figure 1: Procedure planning in instructional videos using diffusion models. Upper section: Procedure planning task is to generate intermediate actions given the start and goal observation. Lower left section: Previous work (Projected Diffusion) [29] does not take the temporal dependencies between actions into account. Lower right section: Our method incorporates these dependencies into the diffusion model.
@@
Diffusion models have achieved outstanding results in many research fields such as image generation [11, 21], text-to-image generation [22, 33, 23], trajectory planning [13, 12], video generation [10, 18], human motion prediction [31], time series imputation and generation [26, 14] and so on. The current state-of-the-art method for procedure planning in videos [29] is also based on a diffusion model. Unlike the diffusion models for images, the input for the diffusion model is a multi-dimensional matrix that consists of the visual observations of the start and goal, the sequence of actions, and task classes. At inference time, the action plan is taken from the action sequence in the generated full matrix. A key limitation of this work is that the influence of the temporal dependencies of actions, i.e. that actions are more likely to cause particular follow-up actions, has not been considered. Although the input matrix in [29] contains the action labels during the training, the diffusion model still treats the input matrix as a static “image”, it does not learn these temporal dependencies.
@@
To overcome this limitation we propose Action-aware Noise Mask Diffusion (ActionDiffusion) – the first method that leverages the temporal dependencies between actions into the diffusion process (see Figure 1). Our method learns the temporal dependencies of actions in the noise space of the diffusion instead of the feature space, thereby unifying the tasks of learning temporal dependencies and generating action plans in the diffusion steps. More specifically, we first propose an action-aware noise mask for the noise-adding stage of the diffusion model. We add the action embeddings in addition to the Gaussian noise in the noise masks so that the model input is transformed into Gaussian noises by iteratively adding the action-aware noise. Second, we introduce an attention mechanism in the denoising neural network (U-Net) to learn the correlations between actions. During the inference, it can then predict action-aware noises to generate action plans. In line with previous work [34, 29], we evaluate the performance of ActionDiffusion through experiments on three popular instructional video benchmarks: CrossTask [35], Coin [25], and NIV [1] with various time horizons. Our results show that ActionDiffusion achieves State-Of-The-Art (SOTA) performances across different metrics for all time horizons on all three datasets. We demonstrate that by adding action embeddings into the noise mask, the diffusion model can effectively learn the temporal dependencies between actions and increase the performance on procedure planning in instructional videos.
@@
2 Related Work
2.1 Learning Actions from Videos
There are several lines of research in learning actions from videos. Action recognition [30, 32] is to classify what actions humans are performing in videos. This is a video classification problem. Action anticipation [6, 7] is the task of predicting future actions based on the video given to a model. The task of procedure planning differs from action recognition and action anticipation, it plans the action sequence between the given visual input of the start and goal state. There were works to learn and plan actions from visual input [9, 27, 8]. However, these works learnt visual representations from rather simple simulated images. Other works used real-world images to learn action plans [5, 24], however, the environment was still simple and constrained. Understanding the natural real-life scenario in which humans are present to plan actions is still a challenge and it requires models to have the ability to understand complex visual scenes and human actions.
2.2 Procedure Planning
The task of procedure planning was defined by [3]. The authors used CNNs to extract visual features and modeled the dynamics between observations and actions in the feature space using Multi-Layer Perceptrons (MLPs) and Recurrent Neural Networks (RNNs). In [24], the authors followed the same paradigm of modelling the dynamics between actions and observations but used transformers instead. RL was used in [2] to learn the action policy. All these works used a separate planning algorithm (Beam search or Walk Through) during planning. Later works used generative models to sample action plans during the inference stage. In [34], a generative component together with transformers was trained for sampling action plans during inference. The denoising diffusion probabilistic model was used for procedure planning in [29]. The action plan, task class, $o_{S}$ and $o_{g}$ were formed as the input to the diffusion model and the action plan was denoised from random noise.
2.3 Modeling Temporal Dependencies
Previous works used different approaches to model the temporal dependencies between actions in procedure planning. In [3], the authors modeled the temporal dependencies based on Markov Decision Process (MDP), where the action at $t+1$ is based on the current state $x_{t}$ and the current action $a_{t}$. As transformers [28] have shown strong capabilities in sequence modeling, P3IV [34] used the transformer to empirically model the temporal dependencies by putting the start and goal observations as the first and last queries in the transformer and making the intermediate queries (related to action sequences) learnable. PlaTe [24] also used MDP for modeling temporal dependencies, specifically two transformers, i.e. an action transformer and a state transformer were used. During training, all actions and states were used and during inference all past action-state pairs were used. Although we also use self-attention to model the temporal dependencies, which is a key component in transformers, there are two differences. First, we build the correlation between actions by accumulating action embeddings in noise-adding stage in the diffusion model. Second, we add self-attention in the U-Net to enhance the temporal relations during denoising.
3 Method
@@
3.1 Problem Formulation
We adopt the problem formulation of procedure planning used in previous work [3, 29]: Given the visual observation of the start state $o_{s}$ and the goal state $o_{g}$, the task is to predict the intermediate steps, i.e., the action plan $\pi=a_{1:T}$ for a chosen time horizon $T$. The action plan $\pi$ will transform $o_{s}$ to $o_{g}$. More formally, procedure planning can be written as,
@@ 3.1 Problem Formulation
where $\hat{c}$ is the predicted task class of the video (e.g., make coffee) and to complete a task, a sequence of actions needs to be performed. The planning is decomposed into two steps [29]: 1) predicting the task class $\hat{c}$ given the start state $o_{s}$ and the goal state $o_{g}$, 2) inferring the action plan $\pi$ given $o_{s}$, $o_{g}$, and $\hat{c}$ by sampling from the diffusion model.
3.2 Action-aware Noise Mask Diffusion
Figure 2 provides an overview of our proposed ActionDiffusion method. Taking an instructional video as input, the method first extracts the visual features of the start state $o_{s}$, the goal state $o_{g}$, and of action embeddings $a_{e_{1:T}}$. It then uses the task class $c$, one-hot action class $a_{1:T}$, $o_{s}$, and $o_{g}$ to form the input for the diffusion model. Note that during training, we use the ground truth task class $c$ while the predicted task class $\hat{c}$ is used during inference. To obtain $\hat{c}$ we train a separate task classifier that models $p(\hat{c}\mid o_{s},o_{g})$ in Equation 1. In the noise-adding phase, during training, the noise is added on $a_{1:T}$ in the model input $x_{0}$. For each action, we add all previous as well as the current action embedding in addition to the Gaussian noise. In the denoising phase, during inference, we use a U-Net with attention to predict the action-aware noise to denoise the noisy input $x_{n}$ at step $n$. The action sequence $\hat{a}_{1:T}$ from the reconstructed input $x_{0}$ is the predicted action plan (see below for details).
3.3 Diffusion Model
@@
3.4 Action-aware Noise Mask
@@
at the noise-adding step, the noise is only added to the action dimension, i.e. $a_{1:T}$ and $a_{i}$ is a one-hot vector of a number of $A$ action classes, thus $\mathbf{I}\in\mathbb{R}^{T\times A}$ in Equation (2) and $x_{0}\in\mathbb{R}^{T\times(O+A+C)}$, $O$ and $C$ are the dimension of $o_{s}$ and the dimension of task class $c$. We design an action-aware noise mask $M_{a}$ with multiple previous actions accumulated (MultiAdd) as
@@
where $a_{e_{1:T}}$ are the embeddings of $a_{1:T}$, and $g(\cdot)$ normalises the values of action embeddings to the range of [-1, 1]. $M_{a}$ has the same dimension as $\mathbf{I}$. Since the denoising step clamps the feature to the range of [-1, 1] too, we normalise the action embeddings for more stable training. In addition to the Gaussian noise applied to $x_{0}$, we add the normalised action embeddings to the noise mask. In the temporal direction, we accumulate all previous action embeddings. Our intuition is that the actions will affect the states and this should be reflected in the noise space. At each $i\in[2:T]$ the noise mask knows what the previous actions are and $\mu_{\theta}$ can learn the temporal dependencies of actions in the denoising stage. With the action-aware noise mask, Equation 2 then becomes,
@@
we again apply re-parameterisation and update Equation 3,
$x_{n}=\sqrt{\bar{\alpha}_{n}}x_{0}+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{a},$ (9)
where $\epsilon_{a}\in\mathcal{N}(\mu_{a},\sigma_{a}^{2}\mathbf{I})$.
3.5 Denoising Neural Network
@@
Following [29], we use a U-Net as the noise prediction neural network $\epsilon_{\theta}$. To better learn the temporal dependencies between actions we further propose to incorporate an attention mechanism [4, 22]. Figure 3 shows the architecture of the U-Net with attention. As indicated in Equation 5, the U-Net takes $x_{n}$ and noise schedule time step $n$ as input. $n$ is first passed to a time step block and then is fed to all residual blocks. The attention block takes $x$ as input and extracts Query ($Q$), Key ($K$) and Value ($V$) by a 1D convolutional layer and calculates the standard self-attention.
3.6 Training
@@
$\mathcal{L}=\sum_{n=1}^{N}(\mu_{\theta}(x_{n},n)-x_{0})^{2},$ (10)
where $\mu_{\theta}$ is the denoising neural network U-Net, $x_{n}$, $x_{0}$ and $n$ are the noised input, the input without noise and the timestep for adding noise respectively.
3.7 Inference
During inference, only $o_{s}$, $o_{g}$ and the predicted task class $\hat{c}$ are given and we need to sample the action plan $\pi$. We start with constructing $x_{n}$. Then, the noise prediction neural network $\mu_{\theta}$ predicts the noise iteratively to denoise $x_{n}$ into $x_{0}$. Here, $x_{n}$ is
$x_{n}=\begin{bmatrix}\hat{c}&\hat{c}&...&\hat{c}&\hat{c}\\ \epsilon_{a_{1}}&\epsilon_{a_{2}}&...&\epsilon_{a_{T-1}}&\epsilon_{a_{T}}\\ o_{s}&0&...&0&o_{g}\end{bmatrix},$ (11)
only the action plan $\pi$ is initialised with noise $\epsilon_{a_{i}}$ and $\epsilon_{a_{i}}\in\mathcal{N}(0,\sigma_{a_{i}}^{2}\mathbf{I})$ since the noise is only added to the actions during the noise-adding stage. After $N$ steps of denoising, we take the action sequence $\hat{a}_{1:T}$ as the action plan $\pi$. In other words, all actions in $\pi$ are generated at once. It is worth noting that we use action-aware noise masks in the noise-adding stage, the action embeddings are added to Gaussian noise (Equation 8). Since the action embeddings are approximately normally distributed (see Figure 4), the noised input after $N$ time steps is also approximately normally distributed with a different standard deviation compared to Gaussian noise (see Figure 5). Hence, instead of sampling from $\epsilon\in\mathcal{N}(0,\mathbf{I})$, we sample from $\epsilon_{a_{i}}\in\mathcal{N}(0,\sigma_{a_{i}}^{2}\mathbf{I})$. The comparison of using $\epsilon$ and $\epsilon_{a}$ for inference is shown in supplementary material.
(a) Crosstask
(b) Coin
(c) NIV
Figure 4: Examples of action embedding distributions from the CrossTask (a), Coin (b), and NIV datasets (c).
4 Experiments
4.1 Datasets
@@
4.2 Implementation Details
We use AdamW [17] as the optimizer for all datasets. For the training on the Crosstask dataset, we train with batch size 256 and 120 epochs. In each epoch, the train step is 200. The first 20 epochs are the warm-up stage, and the learning rate increases to $5e^{-4}$ linearly. In the last 30 epochs, the learning rate decays by 0.5 for every 5 epochs. For the Coin dataset, we train 800 epochs with batch size 256 with 200 train steps in each epoch. The warm-up state is 20 epochs where the learning rate increases to $1e^{-4}$ linearly. The learning rate decreases in the last 50 epochs with a decay rate 0.5. For the NIV dataset, we train 130 epochs with batch size 256. The train step in each epoch is 50. The learning rate increases to $1e^{-5}$ for $T=3$ and $3e^{-6}$ for $T=4$ in the first 90 epochs.
We use the same task predictor and adapt the same training strategy from [29] for predicting the task classes. The task predictor is a 4-layer MLP and it takes the image features of $o_{s}$ and $o_{g}$ as input and outputs the predicted task class $\hat{c}$. The accuracy is over 92% on Crosstask, over 78% on Coin and 100% on NIV.
During inference, we sample $x_{n}$ using $\epsilon_{a}$ instead of $\epsilon$. The noised input approximately follows a normal distribution with a different mean and standard deviation other than the Gaussian noise. The mean shifts and the standard deviation are smaller than one (see Figure 5). We calculate the mean and standard deviations of the noised inputs on the training sets of Crosstask, Coin and NIV and use them for sampling in the inference on the test sets. The standard deviation of each action within the time horizon $T$ for all datasets is shown in Table 1.
Table 1: Mean and standard deviations of $\epsilon_{a[1:T]}$ on the training sets of CrossTask, Coin and NIV.
$T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$
$\mu$ $\sigma_{a}^{2}$ $\mu$ $\sigma_{a}^{2}$ $\mu$ $\sigma_{a}^{2}$ $\mu$ $\sigma_{a}^{2}$ $\mu$ $\sigma_{a}^{2}$ $\mu$ $\sigma_{a}^{2}$
CrossTask -0.27 0.09 -0.54 0.13 -0.81 0.16 -1.09 0.18 -1.35 0.21 -1.62 0.22
Coin -0.04 0.59 -0.08 0.68 -0.11 0.72 -0.14 0.72 - - - -
NIV 0.06 0.11 0.12 0.17 0.19 0.20 0.26 0.23 - - - -
4.3 Metrics